# Machine Learning-Driven Insights into Excitonic Effects in 2D Materials

### v2

@@ 1 Introduction
While Liang et al. (2019) pioneered physics-inspired machine learning for band gap and EBE prediction based on Phillips ionicity theory, their reliance on small datasets and ad-hoc structural maps limits the generalizability of their approach. Additionally, their model only covers AmBn materials, whereas the C2DB has been significantly updated over the past five years [8]. In another study, Lin et al. (2023) used features such as the highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) from the C2DB database to predict EBE in 2D materials [9]. Their model achieved a $R^{2}$ value of 0.80 and a mean absolute error (MAE) of 0.21 eV with a gradient boosting (GB) regression model. Other models have the potential to improve this accuracy, as we will demonstrate in this work. While machine learning has been employed to predict EBE, Bayesian Optimization remains unexplored in this context.
@@ 2.2 Features selection
Pearson correlation coefficient quantifies the linear relationship between variables, providing insight into how one property influences another in monolayer materials. In this study, it serves as a critical tool to evaluate the dependencies between different features [17, 18]. A high Pearson coefficient indicates a strong linear association, such as between PBE and HSE06 band gaps, suggesting that trends in simpler PBE calculations can predict more computationally intensive HSE06 results. Similarly, correlations between G0W0 band gaps and other properties reflect how electronic interactions evolve between computational methods. The wrapper method evaluates the predictive power of features by assessing their impact on model performance. Unlike statistical filter methods, it iteratively trains a model with different feature subsets, capturing both linear and nonlinear dependencies [19]. Understanding these relationships helps optimize feature selection for predicting EBE. Based on these methods, the features selected for this study include the layer thickness, layer group number, atomic number, valence electron count, spin, number of atoms and PBE band gap of the structure.
@@ 2.3 Algorithm selection
To predict the exciton binding energy, we evaluated multiple machine learning algorithms, each with distinct advantages. Neural Networks (NNs) excel at capturing complex, nonlinear relationships by adjusting weights through layers of interconnected neurons [20]. Random Forest (RF) and Gradient Boosting (GB) are ensemble tree-based methods [21]: RF combines predictions from multiple decision trees to reduce variance, while GB sequentially builds trees to minimize prediction error. Support Vector Machines (SVM) define optimal hyperplanes to separate data, useful for high-dimensional spaces [22]. Kernel Ridge Regression (KRR) combines ridge regularization with kernel methods to handle nonlinearity [23]. We used the these models as implemented in scikit-learn [24] package. In the following section, we present the results of these ML models in predicting EBE.
@@ 3.1 Predicting Quasi-particle energies
The Random Forest regression model demonstrates strong performance in predicting G0W0 band gaps, which correspond to quasiparticle (QP) energies. The model’s effectiveness is illustrated in Figure 1. The distribution of prediction errors for the G0W0 band gaps demonstrates a strong alignment with actual values from the C2DB database, indicating the model’s high predictive accuracy and reliability. The histogram of errors has a clear peak around zero, as shown in Figure 2, indicating that most of the predictions are very close to the actual band gap values, R2 equals to 0.98, with MAE and RMSE values of 0.20 and 0.33, respectively.
@@ 3.2 Exciton binding Energy
The derived linear regression equation for EBE demonstrates its dependence on the dominating features.
$EBE=0.31\ E_{PBE}-0.15\ Z-0.06\ d+0.11\ n_{ve}+0.94$ (2)
The equation shows that the PBE band gap is a significant contributor to predicting the exciton binding energy, emphasizing its dominant role. The variable ‘Z’ represents the atomic number, ‘d’ stands for the thickness, and ‘$n_{ve}$’ indicates the valence electron count. This model provides a simple and interpretable relationship between the EBE and the most influential features.
@@ 3.3 Bayesian Optimization
Theoretical predictions indicate that MoS2, the most studied TMDC monolayer, exhibits a large EBE ranging from 526 meV to 1 eV [28]. However, the table suggests there are other candidates with EBEs exceeding that of MoS2, which remain underexplored and merit further investigation.

### v3

@@ 1 Introduction
While Liang et al. (2019) pioneered physics-inspired machine learning for band gap and EBE prediction based on Phillips ionicity theory, their reliance on small datasets and ad-hoc structural maps limits the generalizability of their approach. Additionally, their model only covers AmBn materials, whereas the C2DB has been significantly updated over the past five years [8]. In another study, Lin et al. (2023) used features such as the highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) from the C2DB database to predict EBE in 2D materials [9]. Their model achieved a $R^{2}$ value of 0.80 and a mean absolute error (MAE) of 0.21 eV with a gradient boosting (GB) regression model. Other models have the potential to improve this accuracy, as we will demonstrate in this work. While machine learning has been employed to predict EBE, Bayesian Optimization remains unexplored in this context. This combination of scale, methodology, and accuracy marks a substantive advance over prior ML-based EBE studies.
@@ 2.2 Features selection
Pearson correlation coefficient quantifies the linear relationship between variables, providing insight into how one property influences another in monolayer materials. In this study, it serves as a critical tool to evaluate the dependencies between different features [17, 18]. A high Pearson coefficient indicates a strong linear association, such as between PBE and HSE06 band gaps, suggesting that trends in simpler PBE calculations can predict more computationally intensive HSE06 results. Similarly, correlations between G0W0 band gaps and other properties reflect how electronic interactions evolve between computational methods. The wrapper method evaluates the predictive power of features by assessing their impact on model performance. Unlike statistical filter methods, it iteratively trains a model with different feature subsets, capturing both linear and nonlinear dependencies [19]. Understanding these relationships helps optimize feature selection for predicting EBE. Based on these methods, the features selected for this study include the layer thickness, layer group number, atomic number, valence electron count, atomic number, ionic radii difference, and PBE band gap of the monolayer.
@@ 2.3 Algorithm selection
To predict the exciton binding energy, we evaluated multiple machine learning algorithms, each with distinct advantages. Neural Networks (NNs) excel at capturing complex, nonlinear relationships by adjusting weights through layers of interconnected neurons [20]. Random Forest (RF) and Gradient Boosting (GB) are ensemble tree-based methods [21]: RF combines predictions from multiple decision trees to reduce variance, while GB sequentially builds trees to minimize prediction error. Support Vector Machines (SVM) define optimal hyperplanes to separate data, useful for high-dimensional spaces [22]. Kernel Ridge Regression (KRR) combines ridge regularization with kernel methods to handle nonlinearity [23]. We used these models as implemented in scikit-learn [24] package. In the following section, we present the results of these ML models in predicting EBE..
2.4 Workflow
The workflow of this study proceeds in three stages. First, the dataset is constructed using structural and electronic descriptors obtained from the C2DB database. Second, multiple machine learning models are trained and evaluated. Finally, both the G0W0 quasiparticle band gaps and the exciton binding energies are predicted directly from the selected descriptors using the trained models.
@@ 3.1 Predicting Quasi-particle energies
The Random Forest regression model demonstrates strong performance in predicting G0W0 band gaps, which correspond to QP energies. The model’s effectiveness is illustrated in Figure 1. The distribution of prediction errors for the G0W0 band gaps demonstrates a strong alignment with actual values from the C2DB database, indicating the model’s high predictive accuracy and reliability. The histogram of errors has a clear peak around zero, as shown in Figure 2, indicating that most of the predictions are very close to the actual band gap values, R2 equals to 0.98, with MAE and RMSE values of 0.20 and 0.33, respectively.
@@ 3.2 Exciton binding energy
3.3 SHAP Analysis
To interpret the non-linear predictions of exciton binding energy made by the RF model, we employed SHAP (SHapley Additive exPlanations), a game-theoretic approach that assigns each feature an importance value for individual predictions. SHAP values quantify how much each descriptor increases or decreases the predicted EBE, allowing both a global ranking of feature importance and local insight into directionality [25]. Unlike linear regression coefficients, SHAP is model-agnostic and captures the complex non-linear dependencies present in ensemble methods.
Figure 5 presents the SHAP summary plot for the top features in our model. The PBE band gap emerges as the dominant descriptor, with larger band gaps strongly increasing predicted EBE, consistent with reduced dielectric screening in wide-gap systems. Layer thickness also makes a significant contribution, as it correlates with quantum confinement and dielectric screening, both of which strongly influence exciton binding. The mean atomic number (Zmean) ionic radius, and layer group number show more modest but still notable effects, shaping the electronic environment and excitonic interactions.
Figure 5: SHAP (SHapley Additive exPlanations) summary plot for the RF model predicting EBE.
@@ 3.4 Bayesian Optimization
MoS2 0.52 0.55
WS2 0.53 0.52
WTe2 0.46 0.42
MoS2, the most studied TMDC monolayer, is reported to exhibit an exciton binding energy in the range of 0.5–0.8 eV [12, 29], which is consistent with our model’s predicted value (see Table 2). This agreement with literature benchmarks validates the reliability of our approach. At the same time, Table 2 highlights several other monolayers with predicted EBEs comparable to or exceeding that of MoS2, suggesting promising yet underexplored candidates for future investigations.